PMC in the North path got misaligned while working on the input path. We aligned it back again today. PMC locking method summary:
We were able to achieve fringe visibility of 76.16% (measured by VIS program in labs programmable calculator). Also, we were able to get rid of all other modes enough that they were not visible anymore in the oscilloscope.
Keywords for future search: North Cavity Pre Mode Cleaner modecleaner visibility PMC